OpenStudy (anonymous):

If x and y are two different real numbers and rx=ry then r= (a) 0 (b) 1 (c) x/y

jimthompson5910 (jim_thompson5910):

hint: rx = ry rx - ry = 0 r(x - y) = 0 I'll let you finish

OpenStudy (anonymous):

0

OpenStudy (anonymous):

um.....0...

jimthompson5910 (jim_thompson5910):

r(x - y) = 0 r = 0 or x - y = 0 r = 0 or x = y but we're told that x doesn't equal y, so x = y is false

jimthompson5910 (jim_thompson5910):

this means that r = 0 is the only thing that's left to be true